# H-STAR: LLM-driven Hybrid SQL-Text Adaptive Reasoning on Tables

### v2

@@ 3.2 Efficiency Analysis
Analysis. The results shown in Table 4 emphasize the challenges LLMs face when reasoning over longer tables. Our findings align with previous research, revealing a significant decrease in LLM performance as table sizes increase. While other table manipulation methods experience a drop in performance with larger tables, H-STAR maintains consistent performance across all table sizes.
3.3 Error Analysis
H-STAR
H-STAR vs Others
In contrast, H-STAR proves more effective in both table extraction and reasoning, generating correct answers on 43 samples. Additionally, it demonstrates fewer errors in table extraction (4 missed columns; 16 missed rows) and reasoning (27 of the remaining 80 samples, i.e. 34%).
4 Ablation Study
To evaluate the significance of the two primary stages in H-STAR: a) Table extraction and b) Adaptive reasoning, we conducted an ablation study, removing one stage at a time. For the first condition, the final adaptive reasoning stage was substituted with a basic chain-of-thought reasoning process, while maintaining the table extraction stage. Conversely, the second condition retained the adaptive reasoning stage but omitted the table extraction stage, specifically by removing the row extraction and column extraction steps individually and then collectively. A detailed assessment of the contribution of each stage to the overall performance of the method is shown in Table 5. It highlights the importance of each step in the pipeline. Removing any step can result in a performance loss especially adaptive reasoning which causes a 7% drop for both TabFact and WikiTQ datasets. For a detailed error case study and further analysis, refer to the Appendix E.
Method TabFact WikiTQ
H-STAR 86.51 68.62
w/o row extraction 86.17 66.30
w/o column extraction 84.04 67.03
w/o table extraction 83.79 63.58
w/o adaptive reasoning 79.35 61.47
Table 5: Performance on TabFact and WikiTQ datasets.
4.1 Impact of the Hybrid Approach
We perform a quantitative study of the impact of the hybrid approach in H-STAR. Figures 16, 17 (Appendix E) depict how using only one of the approaches results in insufficient outputs for column and row extraction tasks respectively. Furthermore, Table 6 shows that not employing a hybrid approach results in a significant performance loss.
Method TabFact WikiTQ
CoT 79.05 60.43
H-STAR ${}_{\text{text}}$ 79.43 61.47
Text-to-SQL 52.05 42.03
H-STAR ${}_{\text{sql}}$ 58.50 46.09
H-STAR 86.51 68.62
Analysis. SQL methods though efficient with numerical reasoning, are highly sensitive to data variations like irregular formatting and mixed data types, leading to a performance drop. In contrast, textual reasoning is more resilient, offering fuzzy matching and better interpretation of data. By combining textual and symbolic reasoning, H-STAR uses their complementary strengths, achieving superior performance.
4.2 Symbolic vs Semantic Approaches
As shown in Table 7 removing SQL extraction drops the accuracy to 85.22% on TabFact and 64.39% on WikiTQ, removing text extraction causes a drop to 83.74% and 60.31%, respectively. SQL reasoning removal results in 84.48% and 64.76%, but omitting text reasoning causes the largest drop, to 58.70% and 54.35%.
Method TabFact WikiTQ
H-STAR 86.51 68.62
w/o SQL extraction 85.22 64.39
w/o text extraction 83.74 60.31
w/o SQL reasoning 84.48 64.76
w/o text reasoning 58.70 54.35
Table 7: Performance of H-STAR after systematically removing symbolic part and semantic parts from: 1) table extraction (both row and column extraction); 2) adaptive reasoning.
5 Key Findings
6 Related Work
Table reasoning tasks require the ability to reason over unstructured queries and structured or semi-structured tables. Traditional approaches like TAPAS Herzig et al. (2020), TAPEX Liu et al. (2021), TABERT Yin et al. (2020), TURL Deng et al. (2022), PASTA Gu et al. (2022) work on pre-training language models jointly on large-scale tabular and text data to reason in an end-to-end manner. Advancements in LLMs allow them to learn from in-context examples, reducing inference costs. Text-to-SQL Rajkumar et al. (2022) and Program-of-Thought Chen et al. (2023) use symbolic methods to solve table-based tasks via Python/SQL programs. Textual-based reasoning techniques like Table-CoT Chen (2023) and Tab-CoT Ziqi and Lu (2023) extend prompting methods such as zero and few-shot CoT for tabular reasoning.
Decomposing problems into smaller, manageable tasks has proven effective in solving complex reasoning challenges Zhou et al. (2022); Khot et al. (2022). Recent techniques in table reasoning follow this approach, either by breaking tasks into fixed sub-tasks Cheng et al. (2022); Ye et al. (2023); Nahid and Rafiei (2024) or by employing iterative methods Jiang et al. (2023); Zhang et al. (2023); Wang et al. (2023). BINDER Cheng et al. (2022) is an SQL-based approach that modifies SQL statements to include LLM API calls within SQL statements. ALTER Zhang et al. (2024a) augments both the queries along with the table data. DATER Ye et al. (2023) and TabSQLify Nahid and Rafiei (2024) are table decomposition methods that use semantic and symbolic reasoning respectively. Chain-of-Table Wang et al. (2023) uses a textual reasoning approach to update tables iteratively before the final reasoning step. ReAcTable Zhang et al. (2023) extends the ReAct Yao et al. (2023) framework to table reasoning by employing step-by-step reasoning, iteratively generating sub-tables using external tools like SQL and Python. A detailed explanation has been provided in Appendix C. SYNTQA Zhang et al. (2024b) is an ensemble approach employing an answer selection mechanism selecting between text-to-SQL and text-based models. In contrast, H-STAR follows a fixed two-stage table extraction and reasoning pipeline, applying both symbolic and semantic approaches complementarily at each stage, which leads to improved performance.
7 Conclusion
Limitations
Our current work has focused primarily on a subset of table reasoning tasks using datasets sourced from Wikipedia. While this has laid a solid foundation, it limits exploration into diverse reasoning tasks such as table manipulation, text-to-table generation, and table augmentation, which could provide valuable insights and enhance our approach’s capabilities. Additionally, our method’s generalizability is confined to Wikipedia-based datasets, restricting its application to other domains that require specific domain knowledge, which our current approach lacks. Extending our approach to different domains may necessitate integrating domain-specific knowledge to ensure effective reasoning.
Furthermore, our evaluation has been limited to relatively straightforward table structures. Handling more complex data representations such as semi-structured tables, hierarchical tables, and relational databases remains unexplored territory. These structures present unique challenges that our current approach may not effectively address. Finally, our study has focused solely on the English language, potentially limiting applicability to languages with different linguistic complexities that we have not accounted for.
Ethics Statement
Acknowledgments
We thank the authors of Tabsqlify and BINDER for sharing their code, and Varun Yerram for his help in setting up the implementation process. This research was supported by the Army Research Office under Grant Number W911NF-20-1-0080 and partially funded by ONR Contract N00014-19-1-2620. Lastly, we appreciate the reviewing team’s insightful comments.
Appendix A Supplementary Analysis
A.1 Comparison With Diverse Methods
Tables 8 and 9 show results of other baseline methods on WikiTQ and TabFact datasets respectively. Our method, H-STAR, outperformed all other models by substantial margins.
Model Method Accuracy
Pre-trained/fine-tuned TaPas 48.8
GraPPa 52.7
LEVER 62.9
ITR 63.4
SYNTQA(GPT)* 70.4
TableCoT 48.8
DATER 65.9
BINDER 61.9
ReAcTable 65.8
GPT-3.5-Turbo ReAcTable 52.5
TableCoT 52.4
StructGPT 52.2
ALTER 67.4
SYNTQA(GPT)* 65.2
GPT-3.5-Turbo H-STAR 69.6
Table 8: Comparison of H-STAR with additional methods on WikiTQ dataset. * SYNTQA (GPT) uses GPT-3.5-Turbo to choose between SQL/text reasoning and uses: 1) fine-tuned models for reasoning; 2) GPT-3.5-Turbo for reasoning.
Model Method Accuracy
Pre-trained/fine-tuned Table-BERT 68.1
LogicFactChecker 74.3
SAT 75.5
TaPas 83.9
TAPEX 85.9
SaMoE 86.7
PASTA 90.8
Codex TableCoT 72.6
DATER 85.6
BINDER 85.1
ReAcTable 83.1
GPT-3.5-Turbo ReAcTable 73.1
TableCoT 73.1
ALTER 84.3
GPT-3.5-Turbo H-STAR 85.0
A.2 FeTaQA
Represented in Table 10 are the results for the FeTaQA dataset using the ROUGE scores Lin (2004). We observe an incremental improvement in the scores when compared to other methods on the GPT-3.5-Turbo. The ROUGE scores focus on lexical similarities while ignoring the semantic similarity between predicted and gold outputs. These metrics are known to be insensitive to capturing improvements when using in-context learning since the model does not learn the expected style of the long-form text just from an instruction or a few examples.
Prompting ROUGE-1 ROUGE-2 ROUGE-L
PaLM-2*
DATER 0.63 0.41 0.53
Chain-of-Table 0.66 0.44 0.56
GPT-3.5-Turbo
Table-CoT 0.62 0.39 0.51
TabSQLify 0.58 0.35 0.48
H-STAR 0.62 0.39 0.52
Human Evaluation
We further evaluate the quality of our generations using human evaluation. We refer to Nan et al. (2022) and evaluate our generations on four factors: 1) Fluency based on the nature and grammar of the output; 2) Correctness measuring the degree of correctness of the prediction; 3) Adequacy to measure if all the necessary information is contained in the output; 4) Faithfulness to check whether the generation is grounded to the final extracted table used for reasoning. We ask five internal human annotators to assign a score of 1-5 for each criterion and report the percentage of samples having values of 4 or 5. Table 11 shows the results by five annotators for our method and compares it with other prior methods. The performance pattern of H-STAR is similar to human performance with comparable fluency and faithfulness while matching the trend of having higher adequacy over correctness. While we marginally outperform other methods, there is still scope for improvement.
Method Fluent Correct Adequate Faithful
T5-large 94.6 54.8 50.4 50.4
Human 95 92.4 95.6 95.6
TableCoT 96 82 75 87
TabSQLify 97 88 84 93
H-STAR 96.6 87.6 89.6 94
Table 11: Human evaluation results on FeTaQA dataset.
Qualitative Analysis
In this section, we perform a qualitative analysis on 100 samples of FeTaQA with TabSQLify Nahid and Rafiei (2024). The study aims to compare the generations of H-STAR with TabSQLify by evaluating the answer quality. The answers are evaluated based on their fluency and adequacy when compared with the gold answers. Table 12 shows that H-STAR has better quality generations in 26% samples, as opposed to only 18% for TabSQLify. ‘Both’ indicates samples where both the algorithms perform on par. This can be attributed to the use of the same LLM. Many failures for TabSQLify are derived from an insufficient table, which further highlights the limitations of using a single-view approach for table extraction.
Method % of samples
Both 56%
TabSQLify 18%
H-STAR 26%
Table 12: Comparison of H-STAR with TabSQLify on 100 samples.
Figure 8 showcases such an example where TabSQLify returns ‘information not provided’ because of incorrect of an extracted table. It can be noted that H-STAR generates an answer comparable with the ground truth for the same. Figure 9 visually depicts the outputs from both methods. It illustrates the evaluation process undertaken to compare the quality of the outputs.
Figure 9: Comparison of the outputs with TabSQLify.
A.3 Analysis: Generated Samples
In Table 13, we study H-STAR by considering the total number of samples generated by the LLMs. BINDER and DATER utilize self-consistency techniques to refine their results, while the Chain-of-Table approach adopts a more iterative sample creation process. Specifically, BINDER generates 50 Neural-SQL samples, implementing self-consistency, while DATER employs self-consistency at each step, resulting in 100 samples. In contrast, the Chain-of-Table method opts for a more resource-efficient strategy, producing 25 samples across three steps: ‘Dynamic Plan’, ‘Generate Args’, and ‘Query’. TabSQLify generates the fewest samples, with one generation each for the table decomposition and query steps. Using single outputs can result in a decrease in accuracy when the model fails to return an answer. Thus, H-STAR incorporates self-consistency measures by generating two outputs at each stage to ensure output validity. Each step, except Query, involves two generations for both SQL and Text, i.e., a total of four steps for column and row retrieval. These additional steps serve as a precaution against abnormal LLM generations. Overall, H-STAR is efficient and only uses 6-10 sample generations.
@@ A.3 Analysis: Generated Samples
A.4 Benefit of using Transposed Table

### v3

@@ 3.2 Efficiency Analysis
Analysis. The results presented in Table 4 highlight the challenges LLMs face when reasoning over longer tables. Our findings confirm previous research, showing a significant decline in LLM performance as table sizes increase. While other table manipulation methods suffer from this scaling issue, H-STAR maintains consistent performance across all table sizes. This success is attributed to H-STAR ’s efficient table extraction process, which reduces the table size by removing irrelevant data that would otherwise act as noise and hinder reasoning.
Resource Efficiency.
In Table 5, we analyze H-STAR by examining the total number of samples generated by LLMs. BINDER and DATER employ self-consistency techniques to refine their results, while Chain-of-Table follows an iterative sampling process. Specifically, BINDER generates 50 Neural-SQL samples using self-consistency, whereas DATER applies self-consistency at each step, producing 100 samples. In contrast, Chain-of-Table adopts a more resource-efficient approach, generating 25 samples across three steps: ‘Dynamic Plan’, ‘Generate Args’, and ‘Query’. TabSQLify generates the fewest samples, with a single generation for both table decomposition and query steps. However, relying on single outputs can reduce accuracy when the model fails to return a valid answer. To mitigate this, H-STAR integrates self-consistency by generating two outputs per stage, ensuring output validity. Each step, except ‘Query’, involves two generations for both SQL and text, totaling four steps for column and row retrieval. These additional steps act as safeguards against unreliable LLM outputs. Overall, H-STAR maintains efficiency, requiring only 6–10 sample generations.
@@
3.3 Error Analysis
H-STAR.
H-STAR vs Others.
Analysis. In contrast, H-STAR proves more effective in both table extraction and reasoning, generating correct answers on 43 samples. Additionally, it demonstrates fewer errors in table extraction (4 missed columns; 16 missed rows) and reasoning (27 of the remaining 80 samples, i.e. 34%).
4 Ablation Study
To evaluate the significance of the two primary stages in H-STAR: (a) Table extraction and (b) Adaptive reasoning, we conducted an ablation study, removing one stage at a time. For the first condition, the final adaptive reasoning stage was substituted with a basic chain-of-thought reasoning process while maintaining the table extraction stage. Conversely, the second condition retained the adaptive reasoning stage but omitted the table extraction stage, specifically removing the row extraction and column extraction steps individually and then collectively. A detailed assessment of the contribution of each stage to the overall performance of the method is shown in Table 6. It highlights the importance of each step in the pipeline. Removing any step can result in a performance loss especially adaptive reasoning which causes a 7% drop for both TabFact and WikiTQ datasets. For a detailed error case study and further analysis, refer to the Appendix E.
Method TabFact WikiTQ
H-STAR 86.51 68.62
w/o row extraction 86.17 66.30
w/o column extraction 84.04 67.03
w/o table extraction 83.79 63.58
w/o adaptive reasoning 79.35 61.47
Table 6: Performance on TabFact and WikiTQ datasets.
4.1 Impact of the Hybrid Approach
We perform a quantitative study of the impact of the hybrid approach in H-STAR. Figures 16 and 17 (Appendix E) show how the use of only one of the approaches results in insufficient outputs for column and row extraction tasks, respectively. Furthermore, Table 7 shows that not employing a hybrid approach results in a significant performance loss.
Method TabFact WikiTQ
CoT 79.05 60.43
H-STAR ${}_{\text{text}}$ 79.43 61.47
Text-to-SQL 52.05 42.03
H-STAR ${}_{\text{sql}}$ 58.50 46.09
H-STAR 86.51 68.62
Analysis. SQL methods though efficient with numerical reasoning, are highly sensitive to data variations such as irregular formatting and mixed data types, leading to a performance drop. In contrast, textual reasoning is more resilient, offering fuzzy matching and better interpretation of data. By combining textual and symbolic reasoning, H-STAR uses their complementary strengths, achieving superior performance.
4.2 Symbolic vs Semantic Approaches
As shown in Table 8 removing SQL extraction drops the accuracy to 85.22% on TabFact and 64.39% on WikiTQ, and removing text extraction causes a drop to 83.74% and 60.31%, respectively. SQL reasoning removal results in 84.48% and 64.76%, but omitting text reasoning causes the largest drop, to 58.70% and 54.35%.
Method TabFact WikiTQ
H-STAR 86.51 68.62
w/o SQL extraction 85.22 64.39
w/o text extraction 83.74 60.31
w/o SQL reasoning 84.48 64.76
w/o text reasoning 58.70 54.35
Table 8: Performance of H-STAR after systematically removing symbolic part and semantic parts from: (1) table extraction (both row and column extraction); (2) adaptive reasoning.
5 Key Findings
6 Related Work
Table reasoning tasks require the ability to reason over unstructured queries and structured or semi-structured tables. Traditional approaches like TAPAS Herzig et al. (2020), TAPEX Liu et al. (2021), TABERT Yin et al. (2020), TURL Deng et al. (2022), PASTA Gu et al. (2022) work on pre-training language models jointly on large-scale tabular and text data to reason in an end-to-end manner. Advancements in LLMs, allow them to learn from in-context examples, reducing inference costs. Text-to-SQL Rajkumar et al. (2022) and Program-of-Thought Chen et al. (2023) use symbolic methods to solve table-based tasks via Python/SQL programs. Textual-based reasoning techniques such as Table-CoT Chen (2023) and Tab-CoT Ziqi and Lu (2023) extend prompting methods such as zero- and few-shot CoT for tabular reasoning.
The decomposition of problems into smaller and manageable tasks has proven effective in solving complex reasoning challenges Zhou et al. (2022); Khot et al. (2022). Recent techniques in table reasoning follow this approach, either by breaking tasks into fixed sub-tasks Cheng et al. (2022); Ye et al. (2023); Nahid and Rafiei (2024) or by employing iterative methods Jiang et al. (2023); Zhang et al. (2023); Wang et al. (2023). BINDER Cheng et al. (2022) is an SQL-based approach that modifies SQL statements to include LLM API calls within SQL statements. ALTER Zhang et al. (2024a) augments both the queries along with the table data. DATER Ye et al. (2023) and TabSQLify Nahid and Rafiei (2024) are table decomposition methods that use semantic and symbolic reasoning, respectively. Chain-of-Table Wang et al. (2023) uses a textual reasoning approach to update tables iteratively before the final reasoning step. SYNTQA Zhang et al. (2024b) is an ensemble approach employing an answer selection mechanism that selects between text-to-SQL and text-based models. The works most closely related to our approach are ReAcTable Zhang et al. (2023) and E5 Zhang et al. (2024c). ReAcTable extends the ReAct framework Yao et al. (2023) to table reasoning by employing step-by-step reasoning, where it iteratively generates sub-tables using external tools like SQL and Python. Similarly, E5 introduces a multi-step process for hierarchical table question answering, interpreting table structures, generating and executing code, and performing reasoning over the results to derive answers. In contrast, H-STAR employs a fixed two-stage pipeline for table extraction and reasoning, utilizing both symbolic and semantic approaches complementarily at each stage, resulting in improved performance. A comprehensive comparison of the methodological differences and their implications for performance is provided in Appendix B.
7 Conclusion
Limitations
Our current work has primarily focused on a subset of table reasoning tasks using datasets sourced from Wikipedia. While this has laid a solid foundation, it limits exploration into diverse reasoning tasks such as table manipulation, text-to-table generation, and table augmentation, which could provide valuable insights and enhance our approach’s capabilities. Additionally, our method’s generalizability is confined to Wikipedia-based datasets, restricting its application to other domains that require specific domain knowledge, which our current approach lacks. Extending our approach to different domains may necessitate integrating domain-specific knowledge to ensure effective reasoning.
Furthermore, our evaluation has been limited to relatively straightforward table structures. Handling more complex data representations such as semi-structured tables, hierarchical tables, and relational databases remains unexplored territory. These structures present unique challenges that our current approach may not effectively address. Finally, our study has focused solely on the English language, potentially limiting its applicability to languages with different linguistic complexities that we have not accounted for.
Ethics Statement
Acknowledgments
This research was partly sponsored by the U.S. National Science Foundation (NSF) under Grant No. 2416728, Army Research Office under Grant Number W911NF-20-1-0080, and ONR Contract N00014-23-1-2364. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. We thank the authors of Tabsqlify and BINDER for sharing their code, and Yerram Varun for his help in setting up the implementation process. We also extend our gratitude to Jennifer Sheffield from UPenn for research support. Lastly, we appreciate the reviewing team’s insightful comments.
Appendix A Supplementary Analysis
A.1 FeTaQA
The results for the FeTaQA dataset using the ROUGE scores Lin (2004) are given in Table 9. We observe an incremental improvement in the scores when compared to other methods on the GPT-3.5-Turbo. The ROUGE scores focus on lexical similarities while ignoring the semantic similarity between predicted and gold outputs. These metrics often fail to capture improvements with in-context learning, as the model doesn’t learn the long-form text style from just an instruction or a few examples.
Prompting ROUGE-1 ROUGE-2 ROUGE-L
PaLM-2*
DATER 0.63 0.41 0.53
Chain-of-Table 0.66 0.44 0.56
GPT-3.5-Turbo
Table-CoT 0.62 0.39 0.51
TabSQLify 0.58 0.35 0.48
H-STAR 0.62 0.39 0.52
Qualitative Analysis.
In this section, we perform a qualitative analysis on 100 samples of FeTaQA with TabSQLify Nahid and Rafiei (2024). The study aims to compare the generations of H-STAR with TabSQLify by evaluating the answer quality. The answers are evaluated based on their fluency and adequacy compared to the gold answers. Table 10 shows that H-STAR has better quality generations in 26% samples, as opposed to only 18% for TabSQLify. ‘Both’ indicates samples where both the algorithms perform on par. This can be attributed to the use of the same LLM. Many failures for TabSQLify are derived from an insufficient table, further highlighting the limitations of using a single-view approach for table extraction.
Method % of samples
Both 56%
TabSQLify 18%
H-STAR 26%
Table 10: Comparison of H-STAR with TabSQLify on 100 samples.
Figure 8 presents an example in which TabSQLify returns ‘information not provided’ due to an incorrect table. It can be seen that H-STAR generates an answer comparable with the ground truth for the same. Figure 9 visually depicts the outputs from both methods. It illustrates the evaluation process undertaken to compare the output quality.
Figure 9: Comparison of the outputs with TabSQLify.
Human Evaluation.
We further evaluate the quality of our generations using human evaluation. We refer to Nan et al. (2022) and evaluate our generations on four factors: (1) Fluency based on the nature and grammar of the output; (2) Correctness measuring the degree of correctness of the prediction; (3) Adequacy to measure whether all the necessary information is contained in the output; (4) Faithfulness to check whether the generation is grounded in the final extracted table used for reasoning. We ask five internal human annotators to assign a score of 1-5 for each criterion and report the percentage of samples having values of 4 or 5. Table 11 shows the results by five annotators for our method and compares it with other prior methods. The performance pattern of H-STAR is similar to human performance with comparable fluency and faithfulness while matching the trend of having higher adequacy over correctness. Although we marginally outperform other methods, there is still scope for improvement.
Method Fluent Correct Adequate Faithful
T5-large 94.6 54.8 50.4 50.4
Human 95 92.4 95.6 95.6
TableCoT 96 82 75 87
TabSQLify 97 88 84 93
H-STAR 96.6 87.6 89.6 94
Table 11: Human evaluation results on FeTaQA dataset.
A.2 Comparison With Diverse Methods
Tables 12 and 13 show results of other baseline methods on WikiTQ and TabFact datasets respectively. Our method, H-STAR, outperformed all other models by substantial margins. Despite using the smaller, less capable GPT-4o-mini, H-STAR outperforms E5 and ReAcTable which use GPT-4 on both TabFact and WikiTQ. Furthermore, H-STAR (GPT-3.5-Turbo) outperforms E5 (GPT-4) and ReAct (GPT-4) on WikiTQ.
Model Method Accuracy
Pre-trained/fine-tuned TaPas 48.8
GraPPa 52.7
LEVER 62.9
ITR 63.4
SYNTQA(GPT)* 70.4
TableCoT 48.8
DATER 65.9
BINDER 61.9
ReAcTable 65.8
GPT-3.5-Turbo ReAcTable 52.5
E5 50.9
TableCoT 52.4
StructGPT 52.2
ALTER 67.4
SYNTQA(GPT)* 65.2
GPT-4 ReAcTable 57.3
E5 65.5
GPT-3.5-Turbo H-STAR 69.6
GPT-4o-Mini 74.9
Table 12: Comparison of H-STAR with additional methods on WikiTQ dataset. * SYNTQA (GPT) uses GPT-3.5-Turbo to choose between SQL/text reasoning and uses: (1) fine-tuned models for reasoning; (2) GPT-3.5-Turbo for reasoning.
Model Method Accuracy
Pre-trained/fine-tuned Table-BERT 68.1
LogicFactChecker 74.3
SAT 75.5
TaPas 83.9
TAPEX 85.9
SaMoE 86.7
PASTA 90.8
Codex TableCoT 72.6
DATER 85.6
BINDER 85.1
ReAcTable 83.1
GPT-3.5-Turbo ReAcTable 73.1
E5 75.6
TableCoT 73.1
ALTER 84.3
GPT-4 ReAcTable 83.7
E5 88.8
GPT-3.5-Turbo H-STAR 85.0
GPT-4o-Mini 89.4
A.3 Advantages of using Transposed Table